OpenStudy (anonymous):

What is the solution for x in the proportion 12/22=x/99 ? Describe your steps.

OpenStudy (hoblos):

12/22=x/99 multiply both sides by 99 x= 54

OpenStudy (anonymous):

x=12*99/22 x=1188/22 x=54